5/9+ 5/81+ 2/105=
Simplify the answer into a fraction

Answer by stanbon(48544)

5/9+ 5/81+ 2/105
----------
9 = 3^2
81 =3^4
105=3*5*7

lcd = 3^4*5*7
----------------------

= (5*3^2*5*7)/lcd + (5*5*7)/lcd + (2*3^3*)/lcd
= (1575 + 175 + 54)/lcd
= 1804/2835
===================
Cheers,
Stan H.